OpenStudy (anonymous):

Find the area of a regular octagon with a radius of 10 cm.

OpenStudy (watchmath):

The octagon forms 8 equal triangles each which angle \(360^\circ/8=45^\circ\). Then the area of one of the octagon is \(8\cdot \frac{1}{2}\cdot 10^2\sin 45^\circ=200\sqrt{2}\)

OpenStudy (anonymous):

Thank you! I've had the hardest time finding a formula that makes sense.